OpenStudy (anonymous):

consider the equation x^2 y''-7xy'+15y=0, show that y=x^3(x>0) is a solution of the equation and find a second independent solution.

OpenStudy (anonymous):

hint : this is Euler's homogeneous equation he auxillary equation is D'(D'-1) -7D' +15=0 hence (D'-3)(D'-5)=0 so check if the other solution is of the form y=kx^5 , k being constant ...

OpenStudy (mandre):

x^2 y''-7xy'+15y=0 Using the power rule calculate the 1st and 2nd derivatives: \[y' = 3 x ^{2} \] \[y'' = 6x\] So, substitute for y, y' and y'': \[x ^{2}.(6x) - 7x(3x ^{2}) + 15(x ^{3}) = 0\] \[6x ^{3} - 21x ^{3} + 15x ^{3} = 0\] That's easy math right there. I will defer to matricked for the other solution.